The Scandrick Corporation needs to raise $80 million to finance its expansion into new markets. The company will sell new shares of equity via a general cash offering to raise the needed funds. The offer price is $35 per share and the company’s underwriters charge a spread of 5 percent. If the SEC filing fee and associated administrative expenses of the offering are $600,000, how many shares need to be sold? (Do not round intermediate calculations and enter your answer in shares, not millions of shares, rounded to the nearest whole number, e.g., 1,234,567.)
Answer #1

Required proceeds net of spread=(80,000,000+SEC filing fee and associated administrative expenses)/(1-spread)

=(80,000,000+600,000)/(1-0.05)

=$84842105.3

Hence required shares=84842105.3/35

=2424060 shares(Approx).